OpenStudy (celinegirl):

Few math questions.. I am really behind in math. (: thanks Choose the equation of the horizontal line that passes through the point (1, −5). y = −5 x = 1 y = 1 x = −5

OpenStudy (celinegirl):

If f(x) = 2x2 − 10, find f(5). 90 10 7.5 40

OpenStudy (celinegirl):

Choose the equation of the vertical line passing through the point (−4, 2). y = −4 x = −4 y = 2 x = 2

OpenStudy (celinegirl):

y = x − 3 y = −x + 3 y = −x − 3 y = x + 3

OpenStudy (trojanpoem):

 methods: 1)From the drawing we notice that for any value of x , y = -5 2) equation of staight line y = mx + b slope of horizontal line is 0 because dy/dx = -5 - (-5) / dx = -5 + 5/dx = 0 y = 0 * x + ( y-interception) y = -5

OpenStudy (trojanpoem):

let f(x) = x to find f(x) when x = 2 All you do is replacing each x with the given value , you can even re-write f(x) as y and you have the value of x given. ( let f(x) be y) f( x= 2)= f(2) = 2 , y= x ? what's y ? 2 as x :

OpenStudy (trojanpoem):

To find the equation you can use the slope equation y_2 - y_1/ x_2 - x_1

OpenStudy (trojanpoem):

In both 1, 3

OpenStudy (texaschic101):

1. A horizontal line is represented by y = a number....so the first one is y = -5

OpenStudy (trojanpoem):

vertical line passing through the point (−4, 2). assume another point (x,y) which the line passes through. slope of vertical line = undefined = 1/0 y -2 / x +4 = 1/0 x = -4

OpenStudy (texaschic101):

a vertical line is represented by x = a number....so 3 is x = -4

OpenStudy (trojanpoem):

Last one has a positive slope so y = mx + b ( m > 0) and the interception is 3 so y = x + 3

OpenStudy (texaschic101):

if f(x) = 2x^2 - 10....find f(5) sub in 5 for x f(5) = 2(5^2) - 10 f(5) = 2(25) - 10 f(5) = 50 - 10 f(5) = 40

OpenStudy (celinegirl):

@TrojanPoem @texaschic101 thank you.. sorry for the late reply.

OpenStudy (texaschic101):

its all good :)

OpenStudy (trojanpoem):

Any time.